# Benchmarking XAI Explanations with Human-Aligned Evaluations

### v1

@@ Abstract
In this paper, we introduce PASTA (Perceptual Assessment System for explanaTion of Artificial intelligence), a novel framework for a human-centric evaluation of XAI techniques in computer vision. Our first key contribution is a human evaluation of XAI explanations on four diverse datasets—COCO, Pascal Parts, Cats Dogs Cars, and MonumAI—which constitutes the first large-scale benchmark dataset for XAI, with annotations at both the image and concept levels. This dataset allows for robust evaluation and comparison across various XAI methods. Our second major contribution is a data-based metric for assessing the interpretability of explanations. It mimics human preferences, based on a database of human evaluations of explanations in the PASTA-dataset. With its dataset and metric, the PASTA framework provides consistent and reliable comparisons between XAI techniques, in a way that is scalable but still aligned with human evaluations. Additionally, our benchmark allows for comparisons between explanations across different modalities, an aspect previously unaddressed. Our findings indicate that humans tend to prefer saliency maps over other explanation types. Moreover, we provide evidence that human assessments show a low correlation with existing XAI metrics that are numerically simulated by probing the model.
1 Introduction
As Deep Neural Networks (DNNs) systems are being used in increasingly high stakes domains (e.g., legal, medical) (Surden, 2021; Litjens et al., 2017), it is essential that humans are able to interpret how they reach their conclusions. (Bender et al., 2021). Their lack of transparency has led them to be characterized as “black boxes” (Castelvecchi, 2016), which is particularly problematic in critical applications where understanding the decision-making process is essential for trust and accountability (Vereschak et al., 2024), leading to the creation of a relatively new field: explainable AI (XAI) (Gunning et al., 2019). XAI aims to make the workings of deep learning models more transparent and interpretable. XAI methods fall into two main categories: post-hoc techniques (Selvaraju et al., 2017; Ribeiro et al., 2016; Lundberg & Lee, 2017) and ante-hoc techniques (Bennetot et al., 2022; Koh et al., 2020). Post-hoc techniques generally explain the output of a frozen, pretrained DNN, while ante-hoc techniques modify the architecture of the DNN to improve its interpretability from the outset. Each of these categories can be further subdivided into various sub-families, offering a wide array of XAI approaches.
The diversity of XAI techniques calls for an effort to standardize their evaluation and comparison. Although there are toolkits in computer vision that offer a range of computational evaluation techniques (Hedström et al., 2023; Fel et al., 2022a), to our knowledge there has been no effort to standardize their evaluation from a perceptual point of view (Nauta et al., 2023), i.e., the way the explanation is perceived by the human for whom it was intended. Currently, prevalent approaches (Dawoud et al., 2023; Colin et al., 2022) to evaluating XAI techniques involve human annotators assessing and ranking their interpretability. This approach aligns with XAI’s goal of improving the human interpretability of DNN models. Yet, this method is costly since it requires paying annotators and is impractical for widespread use, as each new XAI technique necessitates a fresh round of human evaluation. It is also at risk of being inconsistent and unreliable since evaluations may differ from one annotator to another and depend on factors such as fatigue and even the time of day (Schmidt et al., 2007).
To address the challenges associated with evaluating XAI techniques, we propose the Perceptual Assessment System for explanaTion of Artificial intelligence (PASTA). PASTA aims to automate the evaluation of XAI techniques by providing an evaluation metric that mimics human assessments.
The first component of PASTA is a dataset composed of four diverse datasets (COCO, Pascal Part, Cats Dogs Cars, and Monumai), which includes both image and concept annotations. Using this dataset, we compare 21 XAI methods across multiple model architectures. We subject the resulting explanations to a rigorous evaluation by human annotators, along a comprehensive set of criteria that cover a variety of desired properties.
The second component of PASTA is a metric designed to replicate human evaluation on the PASTA-dataset. While there are benchmarks that focus on perceptual evaluation of XAI methods (Colin et al., 2022; Dawoud et al., 2023), to the best of our knowledge, we are the first to integrate both saliency-based and concept-based explanations into a unified framework. Additionally, our approach addresses multiple dimensions of human assessment by incorporating a diverse set of questions for users. The primary contributions of this paper are as follows:
  • Comprehensive XAI Benchmark: We establish a dataset, the PASTA-dataset, designed to evaluate XAI methods across various modalities, including image and concept-based explanations (Sect. 3.1).
  • Extensive Evaluation of XAI Methods: We conduct a large-scale evaluation of 21 XAI methods, comparing both post-hoc and ante-hoc approaches across multiple datasets (Sect. 3.23.4). Our findings indicate that saliency and input perturbation-based techniques, such as LIME and SHAP, are favored for their effectiveness in interpreting model predictions (Sect. 3.5).
  • Human-AI Correlation: Our findings reveal a low correlation between widely used XAI metrics and human assessments, suggesting that these metrics cover complementary aspects (Sect 3.6).
  • Human-aligned Perception Metric for Explanations: We introduce a novel, data-based metric, which we call the PASTA-metric, that automates the scoring of XAI techniques along human-like interpretability criteria (Sect. 4).
Automated yet human-aligned metrics such as the PASTA-metric may serve not only to streamline the evaluation process of XAI techniques but also to foster a more transparent and trustworthy AI ecosystem, where deep learning models are comprehensible and their decisions justifiable. The complete PASTA framework (code, annotation, and models) will be released after the reviews.
2 Related Work
@@
Automated scoring involves developing models that assign scores to inputs based on a reference dataset, often derived from human ratings. A particularly active area of research in this domain is automated essay scoring. Traditionally, this has been addressed through handcrafted feature extraction (Yannakoudakis et al., 2011), but modern methods tend to be closer to model as a judge (Lee et al., 2024; Taghipour & Ng, 2016; Chiang et al., 2024). More recently, there has been a growing interest in using embeddings from large language models (LLMs) as features for scoring. The first successful attempt in this direction was made by Yang et al. (2020). Building on this trend, other approaches have incorporated LLM embeddings with models like LSTMs (Wang et al., 2022b), integrated text generation into the training loop (Xiao et al., 2024), or introduced multi-scale aspects to enhance performance (Li et al., 2023a).
Explainable AI.
Explainability refers to computational models designed to provide specific details or reasons to ensure clarity and ease of understanding regarding their functioning (Arrieta et al., 2020). While this issue is not new, there is a growing interest in this field due to the significant role that opaque deep learning models have assumed in various research domains. Traditional machine learning models, such as linear regression or decision trees, inherently offer explanations by providing a computation process simple enough to be understood by humans (Rudin et al., 2022). However, the increasing complexity of state-of-the-art models has necessitated the development of new methods to provide explanations. This has led to the emergence of post-hoc and ante-hoc methods, in this work, we do not propose a new XAI method but a new benchmark that can be used to evaluate new approaches. Additionally, we differentiate between XAI methods, which refer to the processes that generate explanations for various inference models, and XAI techniques, which involve the application of a specific XAI method to a fixed architecture.
Post-hoc methods aim to explain black-box models in human-understandable terms. In computer vision, common techniques include saliency maps (Draelos & Carin, 2020), feature importances (Yuksekgonul et al., 2022), and visual counterfactuals (Augustin et al., 2022). GradCAM (Selvaraju et al., 2017) generates saliency maps from gradient activations, while LIME (Ribeiro et al., 2016) and SHAP (Lundberg & Lee, 2017) use input perturbations to create local model approximations and estimate feature importances, respectively. Selecting fair post-hoc indicators (i.e., that do not mislead interpretation) can be a difficult task (Roy et al., 2022). Ante-hoc explanations offer a compelling alternative, as they incorporate explanations within their design. One approach is to ensure that the latent variables are user-understandable, as seen in Concept Bottleneck Models (CBMs) (Koh et al., 2020), which transform inputs into interpretable concepts before inference. Another method is Chain of Thought reasoning (Ge et al., 2023), which decomposes processing into interpretable subprocesses.
XAI Evaluation.
With the growing interest in XAI, a significant challenge has emerged in comparing existing methods (Lopes et al., 2022). According to Nauta et al. (2023), the absence of standardized benckmarks has limited evaluations to qualitative analyses, often relying on illustrative examples to demonstrate the effectiveness of explanations. Efforts have been made to define key properties for a good explanation, especially regarding audience comprehension. The literature highlights multiple factors influencing explanation quality, leading to various proposed axioms or desiderata, partly due to the task’s subjectivity(Liao et al., 2022). In our work, we consolidate these properties into the following desiderata (more details can be found in Sect. 3.4 and Appendix B.1): faithfulness (Arrieta et al., 2020; Fel et al., 2022b), complexity (Nauta et al., 2023), robustness (Doshi-Velez & Kim, 2017), and objectivity (Bennetot et al., 2022).
Various unsupervised metrics have been developed to provide quantitative values reflecting expected behaviors for specific datasets and methods. For example, Relative Input Stability (Agarwal et al., 2022a) assesses explanation stability under input perturbations, while the deletion metric (Petsiuk et al., 2018) evaluates how removing influential features impacts the model’s prediction. Sparseness (Chalasani et al., 2020), supported by Bénard et al. (2021), argues that good explanations should be concise. Given the range of desiderata for explanations, assessing XAI techniques with a single metric is difficult. Thus, these metrics are often grouped into frameworks like Xplique (Fel et al., 2022a) or Quantus (Hedström et al., 2023), which provide evaluation results across multiple tests.
While automated metrics are useful for evaluating explanations in relation to the model, they inherently overlook the perceptual aspect. As a result, many works involve human annotators who assess methods by selecting preferred explanations for specific examples (Chen et al., 2018; Shu et al., 2019). Notably, significant efforts have been made in the medical field (Miró-Nicolau et al., 2022; Muddamsetty et al., 2021). However, this approach is costly, time-consuming, and hinders reproducibility. To address these issues, benchmarks have been developed to standardize human assessment across various modalities, such as graphical neural networks (Rathee et al., 2022), tabular data (Agarwal et al., 2022b), and language (DeYoung et al., 2020; Wang et al., 2022a). For image processing, solutions for evaluating saliency maps have been proposed by Colin et al. (2022); Mohseni et al. (2021); Dawoud et al. (2023), with Li et al. (2023b) extending this to include textual inputs. However, unlike traditional methods that compare explanations to assumed perfect ground truths, we propose using the model itself as the evaluator. This aligns more closely with human judgment and allows for multimodal explanations, combining concept-based and attribution methods. Moreover, we introduce a multi-value scoring system to better capture users’ diverse expectations.
3 Creating a human preference dataset on XAI interpretability
To evaluate the quality of XAI explanations from a human-centric point of view, we proceed along the following steps, which are detailed in the subsections below. First, we constitute a dataset from annotated images. Using our codebase of classifiers and XAI techniques, we compute label predictions along with XAI explanations. The explanations are then evaluated by human annotators following a rigorous evaluation protocol. Finally, we compare the different XAI techniques with the human evaluations to assess quality of their explanations. We also investigate how human scores correlate with popular automated XAI metrics to see whether they are complementary.
Figure 1: Overview of the human evaluation of the PASTA-dataset. Each image is paired with multiple computed explanations, which are then annotated by human evaluators, first with baseline questions Q0.1 and Q0.2 and then regarding their interpretability and usefulness, using the questions Q1—Q6 outlined in Section 3.4.
3.1 Dataset Composition
To evaluate the performance of different perceptual metrics, we collect a large-scale dataset composed of images of four highly heterogeneous datasets: COCO (Lin et al., 2014), Pascal Part (Chen et al., 2014), Cats Dogs Cars (Kazmierczak et al., 2024), and Monumai (Lamas et al., 2021). This collection is referred to as the classifier’s dataset A.1.1. It is important to distinguish this from the PASTA-dataset, which comprises the final set of images, annotations, labels, and explanations obtained through our evaluation process. Each dataset includes two levels of annotations: image-based annotations and concept-based annotations. This dual-level annotation framework allows for the application of Concept-Based and other XAI methods, enabling a robust evaluation across different approaches. For this, we use for the computation of explanations a subset of 25 images of the classifier’s dataset test split per dataset, resulting in a comprehensive evaluation of 100 images, that serve as the basis of the PASTA-dataset. This diverse selection ensures a broader generalization of the XAI techniques across datasets being assessed. Note that, unlike traditional datasets, our benchmark dataset comprises a triplet of images, explanations, and labels. This triplet enables us to quantitatively assess the quality of XAI techniques (see Figure 1).
3.2 XAI Methods
To ensure representativity, we consider two distinct types of explanations. The first type comprises saliency methods, which generate explanations by assigning an importance score to each pixel of the input image, indicating the significance of each pixel in the prediction process. The second type consists of concept-based explanations, which highlight the importance of human-understandable concepts in the explanation. A detailed list of the methods and more details on each technique are given in Appendix A.2.
Among saliency-based methods, we consider model-agnostic explanations (LIME (Ribeiro et al., 2016) and SHAP (Lundberg & Lee, 2017)), gradient-based (FullGrad (Srinivas & Fleuret, 2019)) and model-specific techniques. The model-specific methods include GradCAM (Selvaraju et al., 2017), HiResCAM (Draelos & Carin, 2020), GradCAMElementWise (Pillai & Pirsiavash, 2021), GradCAM++ (Chattopadhay et al., 2018), XGradCAM (Fu et al., 2020), AblationCAM (Ramaswamy et al., 2020), ScoreCAM (Wang et al., 2020), EigenCAM (Muhammad & Yeasin, 2020), EigenGradCAM (Muhammad & Yeasin, 2020), LayerCAM (Jiang et al., 2021), Deep Feature Factorizations (Collins et al., 2018), and BCos (Böhle et al., 2024).
Among concept-based methods, we explore those that produce explanations through counterfactuals (CLIP-QDA-sample (Kazmierczak et al., 2024)), and feature importance (X-NeSyL (Díaz-Rodríguez et al., 2022), LaBo (Yang et al., 2023), CLIP-linear (Yan et al., 2023), LIME-CBM (Kazmierczak et al., 2024), RISE (Petsiuk et al., 2018) and SHAP-CBM (Kazmierczak et al., 2024)). Additionally, we employ various strategies for concept extraction: zero-shot methods, training from concept annotations, and training from bounding boxes.
3.3 Training Classifiers and Computing the XAI Dataset
The initial phase in constructing the sample set involves training the various classifier models on which explanations will be generated. Specifically, we utilize ResNet50 (He et al., 2016), ViT-B (Dosovitskiy, 2020), ResNet50-BCos (Böhle et al., 2024), CLIP-Linear (Yan et al., 2023), CLIP-QDA (Kazmierczak et al., 2024), X-NeSyL (Díaz-Rodríguez et al., 2022), and ConceptBottleneck (Koh et al., 2020). These models are trained separately on the four datasets. Further details regarding the datasets and training procedures are provided in Section A.1 . The final assessment of XAI techniques is conducted on samples of the test set. To provide a diverse range of explanations, our framework incorporates both black-box models, explained using post-hoc methods, and ante-hoc methods, specifically CBMs. The specific details regarding these two families of methods are presented in Appendix A.1.2.
Our codebase includes the 21 XAI methods described in Sect. 3.2 and the 7 backbone models described above. Some XAI methods are incompatible with certain backbones, see Table 6, this leaves 46 distinct combinations of XAI methods and backbones, which we refer to as XAI techniques. We apply each technique to 100 images. This leads to an XAI dataset of 4600 instances, each of which is associated with an image and its ground truth label, the label prediction from a classifier instance, and the explanation from a particular XAI technique. In the next section, we present an approach for evaluating the human perception of these instances.
3.4 Human evaluation protocol
Figure 2: Samples corresponding to questions 5 and 6. On the left, a light perturbation is applied, resulting in no label change. On the right, a strong perturbation is applied, leading to a label change.
We aim to quantify the interpretability and usefulness of XAI techniques accurately, using a human evaluation of the quality of evaluations. The resulting dataset serves as a benchmark, enabling us to compare and validate current and future XAI methods. Our human-centric approach complements existing approaches that focus primarily on assessing the model’s internal behavior. For example, traditional evaluations of faithfulness measure how closely an explanation corresponds to the model’s true functioning while we assess in our dataset how the explanation fit human expectations.
We take a structured approach to ensure that the explanations are not only technically sound but also align with human cognitive processes and expectations, fostering the development of more transparent and interpretable AI systems. First, we establish a comprehensive set of assessment criteria that are evaluated on a graded scale. Then, we apply a meticulous evaluation protocol, developed with the help of a psychologist, to ensure that annotators fully understand the task and the expectations. This includes annotator training and close monitoring throughout the process.
Evaluation Criteria
We consolidate different criteria from the literature into the following set of desiderata for XAI explanations that we wish to evaluate (for details, see Appendix B.1):
  • Faithfulness (Arrieta et al., 2020; Fel et al., 2022b) measures the extent to which an explanation accurately reflects the true behavior of the model.
  • Robustness (Doshi-Velez & Kim, 2017) assesses the stability and relevance of the explanation across a broad range of models and inputs.
  • Complexity (Nauta et al., 2023) checks whether the explanation is both simple and informative, balancing clarity and detail.
  • Objectivity (Bennetot et al., 2022) evaluates whether the explanation is interpreted consistently by the majority within a given audience.
Evaluation Protocol
Gathering human preferences from surveys is an active field of research in psychology (Fowler Jr, 2013), but also in the field of machine learning e.g., with the recent advent of reinforcement learning from human feedback (RLHF) (Kaufmann et al., 2023). Interfaces, as well as the formulation of questions, play a key role in the quality of the annotations (Pommeranz et al., 2012), and their design must be considered seriously in order to avoid cognitive biases. The following human evaluation protocol has been designed with the help of a psychologist. The formulation of the questions has been carefully chosen to ensure that they are fully understood by each annotator.
To maintain consistency and reliability, all annotators undergo a training session before starting the actual annotation task. This training familiarizes them with the XAI techniques, evaluation criteria, rating scale, and datasets, ensuring a uniform understanding of the task and the expectations. More details about the annotation protocol are given in Appendix A.3, including an example of how questions are presented to the annotators (Fig. 9).
During the evaluation process, annotators are shown an image, a prediction, and an explanation. They are then asked a list of questions that we now describe in more detail. A first set of questions aims at having annotators establish a baseline, i.e., by interpreting and explaining what makes an image recognizable as a specific object or class. This prompts the human annotator to think about what they are relying on to classify the image themselves, before having to evaluate explanations produced by the XAI techniques. The first two questions in the annotation process are:
Each explanation generated by the XAI techniques is then evaluated regarding each of the desiderata given above (indicated in italics):
@@ 3.4 Human evaluation protocol
  • Q1: The provided explanation is consistent with how I would explain the predicted class? Fidelity
  • Q2: Overall the explanation provided for the model prediction can be trusted? Complexity
  • Q3: Is the explanation easy to understand? Complexity
@@
  • Q4: The explanation can be understood by a large number of people, independently of their demographics (age, gender, country, etc.) and culture? Objectivity
Annotators evaluate how well the explanations conform to their expectations (Q1), whether the explanations are clear (Q2 and Q3), and if they would rely on these explanations (Q4). They also assess how much the explanations change when the images are perturbed (Q5 and Q6), for both accurate and inaccurate predictions, as studied by (Fel et al., 2022b). An example is shown in Figure 2. Each explanation is rated on a scale from one to five stars, where one star indicates the explanation is entirely uncorrelated with the annotator’s reasoning, and five stars represent perfect correlation. This star rating system allows for a nuanced assessment of the quality of the explanations, reflecting how closely they align with human understanding.
3.5 Human Evaluation and Results
@@
(a) Scores for each question, for saliency-based and CBM-based explanation
(b) Scores for each question, for different backbones of saliency methods
@@
Figure 3: Comparison of explanation methods and backbones: Overall, saliency methods are preferred over CBM-based explanations. As backbones for saliency methods, ViT-B and CLIP obtain overall similar results, while Resnet50 has better scores.
We now present a brief summary of the human evaluations obtained in the PASTA-dataset, using the previously described protocol. Full results and values are available in Appendices B.2 and B.3. Our XAI dataset, described in Sect. 3.3, contains 4600 instances with images, predictions, and explanations. From this set, we select 2200 samples randomly for the saliency maps as detailed in Appendix A.1 and let them be evaluated by humans according to the protocol above. Each instance receives five evaluations from different annotators. We aggregate these evaluations using majority voting to favor consensus opinions.
As illustrated by Figure 3(a), we observe that these results indicate a preference for image-based techniques, suggesting that saliency maps are perceived as more interpretable than CBMs. Additionally, Figure 3(b) shows the average score among techniques that share the same backbone. Interestingly, CLIP and ViT have similar scores, likely due to the architectural similarities between the two models. ResNet 50, which played a pivotal role in the development of many XAI methods, consistently scores higher. This could suggest a potential bias toward ResNet 50 in the design and effectiveness of current XAI methods.
3.6 Correlation with other metrics.
We turn to the question of how the human scores in the PASTA-dataset correlate with standard XAI metrics. An analysis based on the Pearson Correlation Coefficient and the Spearman rank Correlation Coefficient for different perturbation strategies, shown in Table 1 indicates a rather weak correlation between human scores and ROAD (Rong et al., 2022), a popular metric to evaluate faithfulness. We conclude that our human scores indeed cover an aspect of explanation quality unrelated to that of perceptual quality, as predicted by Biessmann & Refiano (2021). Additional results, including results for other axioms, are available in Appendix B.1 and B.2.
Table 1: Pearson Correlation Coefficient (PCC) and Spearman rank Correlation Coefficient (SCC) between faithfulness computed with different perturbation strategies and human scores. In parentheses, the respective p-values.
ROAD Black patches Uniform noise Gaussian noise
PCC SCC PCC SCC PCC SCC PCC SCC
Q1 0.01 (0.62) -0.04 (0.12) 0.06 (0.02) 0.05 (0.04) 0.07 (0.01) 0.03 (0.24) 0.07 (0.02) 0.02 (0.38)
Q2 -0.01 (0.88) -0.03 (0.20) 0.04 (0.19) 0.04 (0.17) 0.06 (0.04) 0.02 (0.45) 0.05 (0.08) 0.01 ( 0.63)
Q3 0.03 (0.33) -0.03 (0.34) 0.04 (0.10) 0.04 (0.16) 0.08 (0.01) 0.04 (0.13) 0.06 (0.02) 0.03 (0.30)
Q4 0.03 (0.25) -0.04 (0.17) 0.04 (0.19) 0.03 (0.33) 0.08 (0.01) 0.03 (0.21) 0.06 (0.02) 0.02 (0.41)
Q5 -0.05 (0.05) 0.02 (0.41) 0.01 (0.93) 0.04 (0.14) -0.05 (0.07) -0.01 (0.78) -0.04 (0.16) 0.01 (0.95)
Q6 -0.02 (0.37) 0.06 (0.03) -0.13 (1e-5) -0.08 (0.01) -0.04 (0.16) -0.01 (0.68) -0.04 (0.13) -0.01 (0.79)
4 Developing a Metric for Perceptual Evaluation

### v2

@@ Abstract
We introduce PASTA (Perceptual Assessment System for explanaTion of Artificial Intelligence), a novel human-centric framework for evaluating eXplainable AI (XAI) techniques in computer vision. Our first contribution is the creation of the PASTA-dataset, the first large-scale benchmark that spans a diverse set of models and both saliency-based and concept-based explanation methods. This dataset enables robust, comparative analysis of XAI techniques based on human judgment. Our second contribution is an automated, data-driven benchmark that predicts human preferences using the PASTA-dataset. This scoring called PASTA-score offers scalable, reliable, and consistent evaluation aligned with human perception. Additionally, our benchmark allows for comparisons between explanations across different modalities, an aspect previously unaddressed. We then propose to apply our scoring method to probe the interpretability of existing models and to build more human-interpretable XAI methods.
Figure 1: PASTA automates the evaluation of human perception of provided explanations by computing a PASTA-score. By integrating PASTA-score (y-axis) with existing faithfulness metrics (x-axis), we aim to foster the development of explanations that are not only aligned with the model’s behavior but also comprehensible to human evaluators. Samples reported in the figure correspond to the label dog for a ResNet50 classifier trained on PascalPART. XAI methods: top left: GradCAM; Top right: FullGrad; Bottom left: SHAP; Bottom right: AblationCAM
Introduction
As Deep Neural Networks (DNNs) are increasingly deployed in high-stakes domains such as law and medicine (Surden 2021; Litjens et al. 2017), understanding their decision-making process has become essential (Bender et al. 2021). Their opacity often earns them the label “black boxes” (Castelvecchi 2016), raising trust and accountability concerns in critical applications (Vereschak et al. 2024). This has given rise to the field of eXplainable AI (XAI) (Gunning et al. 2019).
A wide variety of XAI techniques have been proposed (Speith 2022; Saeed and Omlin 2023), notably saliency-based methods (Muhammad and Yeasin 2020; Böhle et al. 2024), which highlight relevant input features, and concept-based methods (Yan et al. 2023; D´ıaz-Rodr´ıguez et al. 2022), which associate predictions with high-level semantic concepts. However, comparing such heterogeneous approaches remains an open problem.
Evaluating XAI methods is particularly challenging for two main reasons. First, the diversity of explanation types complicates the definition of a common evaluation framework. Second, the notion of a “good explanation” is inherently subjective. This creates a dichotomy between non-perceptual evaluations—focused on model-centric metrics using toolkits like OpenXAI, Quantus, and Xplique (Agarwal et al. 2022b; Hedström et al. 2023; Fel et al. 2022a)—and perceptual evaluations, which assess human understanding. While the latter is often explored through anecdotal examples (Selvaraju et al. 2017; Wang et al. 2020), user studies (Dawoud et al. 2023; Colin et al. 2022), or region-of-interest alignment (Liu et al. 2024a; Arras, Osman, and Samek 2022), there is still a lack of standardization in assessing explanations from the human perspective (Nauta et al. 2023). Yet, this dimension is crucial—explanations faithful to the model’s reasoning may still be unintelligible to human users, limiting their actionability. In this sense, our proposed PASTA-score allows for evaluating explanations based on a combination of faithfulness and human preferences, as shown in Figure 1.
To address these challenges, we propose PASTA—the Perceptual Assessment System for explanaTion of Artificial intelligence—which aims to automate the human-aligned evaluation of XAI methods. PASTA has two core components. First, a benchmark, PASTA-dataset, composed of four diverse image-based classification datasets with aligned concept annotations, enabling the comparison of 20 XAI techniques across multiple architectures. Second, the PASTA-score, a data-driven metric designed to predict human preferences, providing a scalable way to evaluate explanations from a perceptual standpoint. Unlike prior benchmarks focused solely on saliency or user studies (Colin et al. 2022; Dawoud et al. 2023), PASTA unifies both saliency-based and concept-based methods under a single evaluation framework.
Our contributions are: (1) Comprehensive XAI Benchmark: We introduce the PASTA-dataset, enabling the evaluation of both visual and concept-based explanations. (2) Large-scale Method Evaluation: We assess 20 XAI methods across multiple datasets and models, including both post-hoc and ante-hoc methods. Our first result suggests that human annotators tend to prefer saliency and perturbation-based techniques, like LIME and SHAP. (3) Human-aligned Explanation Scoring: We propose PASTA-score, an automated yet perceptually grounded assessment of explanations, trained on the PASTA-dataset to replicate human preferences. (4) Practical Applications: We present three use cases of PASTA-score, showing how it can guide the design of more interpretable models, and serve as a proxy for visual human assessment in practical deployments. The pipeline of the global workflow is presented in Figure 2.
The complete PASTA framework (code, annotation, and models) will be released upon acceptance.
Figure 2: Pipeline of the PASTA framework. We first collect a dataset of human preferences called PASTA-dataset. This dataset is used to learn to emulate human preferences for the new samples on the test set using the PASTA-score. The trained PASTA-score can then be deployed to downstream applications as a consistent-over-time, quick and cost-effective replacement for human feedback.
Related Work
Explainable AI.
To address the challenge of explaining DNNs, several specialized tools have been proposed, often categorized into post-hoc and ante-hoc methods (Arrieta et al. 2020; Rudin et al. 2022). Post-hoc methods encompass any tool external to the model, allowing us to gain insights from any pre-trained DNN. Popular examples are GradCAM (Selvaraju et al. 2017), LIME (Ribeiro, Singh, and Guestrin 2016), and SHAP (Lundberg and Lee 2017). While most post-hoc explainers agree in providing input regions most responsible for a certain prediction, they differ in many non-trivial details, and selecting and evaluating the most appropriate explainer for each task can be challenging (Leavitt and Morcos 2020; Roy et al. 2022). Ante-hoc methods, instead, aim at modifying the underlying model architecture to provide explanations by design. This can be done in the framework of Concept Bottleneck Models (CBMs) (Koh et al. 2020) by prompting the model to first predict a set of human-understandable high-level concepts, and then making the final prediction using a shallow and interpretable classifier that supports human inspection, or by decomposing the reasoning of the model into smaller and more actionable steps (Ge et al. 2023).
Evaluating explainability.
While several methods have been proposed to quantitatively measure explanation quality, such as faithfulness (Petsiuk, Das, and Saenko 2018; Dasgupta, Frost, and Moshkovitz 2022; Azzolin et al. 2025), sparsity (Chalasani et al. 2020; Bénard et al. 2021), robustness (Alvarez-Melis and Jaakkola 2018b; Montavon, Samek, and Müller 2018), sensitivity (Adebayo et al. 2018; Hedström et al. 2024) and alignment to an assumed ground truth (Colin et al. 2022; Mohseni, Block, and Ragan 2021; Dawoud et al. 2023), they inherently overlook the perceptual aspect with respect to the human, who is the expected consumer of such explanations. Evaluating explanations via user studies, e.g., where annotators are asked to rate and evaluate explanations (Chen et al. 2018; Shu et al. 2019; Yang et al. 2022; Kares et al. 2025), are, however, very costly, prone to unreproducibility issues (Nauta et al. 2023), and often unfeasible for tasks that require trained users, like in the medical domain (Miró-Nicolau, Moyà-Alcover, and Jaume-i Capó 2022; Muddamsetty, Jahromi, and Moeslund 2021). In this work, we take the first step towards standardizing the evaluation of human perception preferences of explanations (Nauta et al. 2023). We propose to overcome the issues of hard-to-reproduce large-scale user studies by automating the evaluation of XAI techniques through a multi-value scoring method that mimics human preferences while taking into account the users’ diverse expectations, which naturally emerge in user-based studies.
@@ Introduction
Automated scoring involves developing models that assign scores to inputs based on a reference dataset, often derived from human ratings. A particularly active area of research in this domain is automated essay scoring. Traditionally, this has been addressed through handcrafted feature extraction (Yannakoudakis, Briscoe, and Medlock 2011), but modern methods tend to be closer to model as a judge (Lee et al. 2024; Taghipour and Ng 2016; Chiang et al. 2024). More recently, there has been a growing interest in using embeddings from large language models (LLMs) as features for scoring. The first successful attempt in this direction was made by Yang et al. (2020). Building on this trend, other approaches have incorporated LLM embeddings with models like LSTMs (Wang et al. 2022), integrated text generation into the training loop (Xiao et al. 2024), or introduced multi-scale aspects to enhance performance (Li et al. 2023).
Creating the PASTA-dataset
To assess the quality of XAI explanations for image classification decisions from a human-centric perspective, we constructed a comprehensive dataset comprising images, predictions, explanations, and evaluations of these explanations, as depicted in Figure 3. To account for the heterogeneity of different XAI methods, model backbones, and training scenarios, we constructed the PASTA-benchmark to include 1000 images sampled across 4 available datasets, 7 classification backbones, 20 XAI methods, 6 questions, and 5 annotations per image. The challenge in annotating such a dataset resides in its multiplicative nature, where each question requires an annotation across multiple backbones, datasets, XAI methods, images, and human annotators. Consequently, the PASTA-dataset contains an overall number of 633,000 samples, each corresponding to a unique Likert-like rating, which is the largest benchmark available of this kind.
To construct such a dataset, the initial phase involved developing a unified platform designed to integrate various models, XAI methods, and datasets in a streamlined manner. This platform encompasses a diverse array of models and XAI methods. Given the potential utility of this platform as a baseline for future research, we intend to release it publicly upon publication of this paper. Further details regarding the overall procedure, including details about the datasets employed, model training, and explanation extraction, are available in Section A.1 of the appendix.
Figure 3: Overview of the human annotation process in the PASTA-dataset. We compute a total number of 46 explanations for each image, out of which 21 are sampled and rated by humans according to six questions. Further details available in Appendix A.2.
The subsequent phase in dataset creation is dedicated to annotating the explanations, which involved 24 participants in an online process. A pivotal insight from existing user study literature (Xuan et al. 2025; Liao et al. 2022) is that human perception of explanations is not unidimensional; rather, it encompasses a range of potentially unaligned desiderata. For instance, a saliency-based explanation that highlights a dog to predict a cat can be entirely clear, thereby satisfying complexity desiderata, while simultaneously not fulfilling plausibility desiderata. Consequently, we pose multiple questions designed to encompass a spectrum of human assessment as broad as possible. Specifically, the following questions were asked:
@@
  • Q1: Is the provided explanation consistent with how I would explain the predicted class?
  • Q2: Overall the explanation provided for the model prediction can be trusted?
  • Q3: Is the explanation easy to understand?
@@
  • Q4: Can the explanation be understood by a large number of people, independently of their demographics (age, gender, country, etc.) and culture?
  • Q5: With this perturbed image, to what extent has the explanation changed ? (Examples with good predictions and light perturbations)
  • Q6: With this perturbed image, to what extent has the explanation changed? (Examples with bad predictions and strong perturbations)
The selection of these questions reflects a broader discourse on user studies and desiderata. To ensure that annotators comprehensively understand the task and expectations, we implemented an evaluation protocol developed with the assistance of a psychologist. This protocol includes annotator training and continuous monitoring throughout the process. Discussions regarding the desiderata, a detailed evaluation protocol, and information about the annotators are available in Section A.2 of the Appendix.
Analysis of Human Preference
@@
Figure 4: Scores for each question, for saliency-based and CBM-based explanation. Overall, saliency-based explanations are preferred over CBM-based ones.
@@
Figure 5: Scores for each question, for different backbones of saliency methods. As backbones for saliency methods, ViT-B and CLIP obtain overall similar results, while ResNet50 has generally better scores.
Having collected a large number of human preferences for different XAI models and backbones in the PASTA-dataset, we now proceed to analyze human scores in relation to each method. The full analysis is available in Appendix B.3.
Human preferences for output format: As illustrated in Figure 4, results indicate that humans tend to prefer image-based explainers in relation to questions Q1-Q5, meaning that saliency maps are perceived as more interpretable than concept-based explanations. Although several factors may contribute to this behavior—such as the lower cognitive load of image-based explanations compared to concept-based ones—a comprehensive investigation of the underlying psychological causes is left for future work. The sole exception to this observation pertains to Q6 (note that for Q5, the ratings are inverted, with a low score indicating favorable behavior). This phenomenon can be explained by the fact that presenting explanations as a heatmap overlaid on the image reduces the perceptual impact of perturbations.
Human preferences for model architecture: Figure 5 illustrates the average scores across XAI methods that use the same backbone, highlighting a general preference for ResNet50. CLIP and ViT achieve similar scores, likely due to the architectural similarities between the two models. ResNet50, which played a pivotal role in the development of many XAI methods, consistently scores higher. This could suggest a potential bias toward ResNet50 in the design and effectiveness of current XAI methods. The results for the last two questions may be due to ViT being more sensitive to the perturbations used than Resnet50. Among the methods we studied, those based on feature factorization—like EigenCAM (Muhammad and Yeasin 2020) and Deep Feature Factorization (Collins, Achanta, and Susstrunk 2018)—tend to give more consistent and preferred explanations. This may be because they remove complex components that can create confusing artifacts, making the explanations easier to understand.
Developing the PASTA-score